# Foundation Models in Electrocardiogram: A Review

### v1

@@ 1 Introduction
In contrast, ECG foundation models (ECG-FMs) present a more versatile solution, offering several advantages [97]. On the one hand, they can perform multiple tasks, such as arrhythmia detection, HRV analysis, and even more complex tasks like predicting heart failure or stratifying patients based on risk profiles [118]. For example, FMs trained on large ECG datasets, like the PTB-XL [103] or PhysioNet’s MIMIC-III [44] waveform database, have been shown to generalize across multiple tasks with high accuracy. These models leverage the diversity of the data, learning features that are transferable across tasks and patient populations, making them better suited for real-world clinical applications where data variability is common. On the other hand, ECG foundation models can incorporate multi-modal data, integrating ECG signals with other physiological or clinical data such as photoplethysmography (PPG) or electronic health records (EHRs), to provide a more comprehensive picture of patient health [86, 3].
@@
In contrast to previous surveys which concentrated on general-purpose FMs or FMs covering a wide scope of medical fields, this article focuses on the ECG-based diagnosis domain. We summarize the developmental roadmap of current ECG-FMs in Figure 1, to inspire advancements and deepen comprehension in the continuously evolving landscape of FMs for ECG analysis. In brief, our major contributions are summarized as follows.
  • Comprehensive and up-to-date review. We thoroughly review the existing literature on ECG-FMs, and provide a comprehensive and up-to-date review on FMs for ECG analysis, encompassing ECG-related datasets and benchmarks, along with a wide spectrum of diagnostic tasks for cardiac diseases.
@@
The rest of the article is organized as follows. Section 2 describes the brief background of FMs. Section 3 presents the methodology of ECG-FMs. Section 4 discusses the existing challenges and promising directions of ECG-FMs, and Section 5 concludes the article.
@@ 2 Background of Foundation Models
FMs are large-scale deep-learning models trained on diverse and extensive datasets in a self-supervised or semi-supervised manner [12]. FMs generally follow the pre-training and fine-tuning paradigm, which enables the model to learn a broad range of knowledge and patterns and to leverage such prior information in various downstream tasks and applications. FMs only require considerably small-scale custom data in the fine-tuning phase, showcasing remarkable adaptability and efficiency [117]. FMs have been adopted successfully across domains such as CV, NLP, and healthcare. As the data or model size massively scales up and the available data modality dramatically increases, FMs have evolved from the initial pre-trained models like BERT to contemporary large language models (LLMs), large vision models (LVMs), vision-language models (VLMs), and time-series foundation models (TSFMs).
@@
Large Vision Models.
Inspired by the success of FMs in NLP, practitioners have delved into the realm of vision FMs, encompassing LVMs and VLMs. LVMs, pre-trained on extensive image datasets, excel in comprehending image content and extracting rich semantic information. With the escalating data scale and the growing complexity of deep neural networks, LVMs have evolved from conventional CNN-based models, such as VGG [100] and EfficientNet [106], to cutting-edge Transformer-based architectures, including ViT [20], Swin Transformer [72], Linformer [114], and DeiT [108]. Moreover, many LVMs employ SSL techniques to acquire generalizable visual representations during pre-training. A prime example is MAE [38], which focuses on the correlation between visible patches and masked regions, enabling it to handle intricate visual structures and capture high-level image semantics.
@@
In addition to LVMs, considerable research efforts have also been channeled into the development of VLMs. VLMs typically incorporate a combined encoder-decoder architecture, leveraging an LLM as the core element, alongside a visual encoder used to transform images into a format compatible with the LLM. For instance, CLIP [89] utilizes contrastive learning methods to create fused representations for images and texts, demonstrating promising capabilities in open-vocabulary image classification. Similarly, SAM [48] stands out as a VLM that delivers impressive zero-shot image segmentation performance, enabling class-agnostic segmentation based on the image and a visual prompt specifying the targeted segments. These pre-trained VLMs can then be tailored to specific tasks by presenting them with a natural language description of the task and corresponding prompts. Recent studies have showcased the promising zero-shot performance of pre-trained VLMs across various downstream vision tasks, such as image classification, object detection, and semantic segmentation [73]. Additionally, research endeavors have also explored the development of models aligning multiple paired modalities, including image-text, video-text, and video-audio [35, 30].
@@
The increasing richness of time-series data and its wide range of applications have thrust TSFMs into the spotlight as a promising direction for time series analysis [43, 60]. TSFMs aim to harness the power of the FM paradigm to develop generalized models adept at understanding and forecasting time series data across diverse domains. By leveraging large-scale time series datasets, TSFMs hold the promise of attaining superior performance across a spectrum of time series tasks, offering a unified framework that can expedite research and application development in this field. Based on the nature of time series, TSFMs can be classified into general-purpose TSFMs dealing with standard time series and domain-specific TSFMs focusing on particular tasks such as traffic forecasting and clinical diagnosis. General-purpose TSFMs predominantly utilize single or multiple data modalities to craft robust models targeting specific time series tasks, such as Lag-Llama [94], TimeGPT-1 [28], TTMs [23], and TSMixer [22] for zero-shot forecasting, and Voice2Series [121], TS2Vec [129], and SimMTM [19] for classification. Domain-specific TSFMs have distinct focuses and architectures tailored to the requirements of the downstream tasks. In the transportation sector, TSFMs typically employ graph structures, such as GNNs, to capture the spatio-temporal interactions within dynamic traffic systems. For instance, STGCL [70] and SPGCL [59] explore the integration of contrastive learning into spatio-temporal graph forecasting, while ST-LLM [63] combines spatio-temporal information with a partially frozen LLM to enhance traffic predictions. Additionally, TSFMs in the medical domain usually place greater emphasis on capturing the historical and hierarchical characteristics of patient data. For example, COMET [115] utilizes a multi-granularity contrastive learning framework to exploit data consistencies across different levels (e.g., observation-level and patient-level) in various medical time series, showcasing its potential benefits in complex disease diagnosis.
@@ 3.1 Architecture Design
As shown in Figure 2, we first delve into the architecture of ECG-FMs, including backbone models and Transformer structure, concentrating on the underlying mechanisms that determine their capabilities and the applications across different ECG-based scenarios.
@@
The original Transformer model utilizes an encoder-decoder architecture with a stack of multi-head self-attention blocks. However, the most suitable transformer model for ECG analysis is still a subject of debate, with options including encoder-only, decoder-only, and encoder-decoder ECG-FMs. Each model offers distinct properties and is suitable for various downstream tasks. The encoder-only structure processes the entire input ECG sequence at once and predicts based on the understanding of the full context, making it ideal for tasks that require interpreting the entire input data, such as cardiac disease detection [122]. On the other hand, the decoder-only architecture demonstrates strong generalization capacity for large-scale ECG pre-training and is well-suited for zero-shot inference [62, 127], ECG synthesis [16], and ECG chatbot [64, 128] applications. Additionally, prominent research combines an encoder-only model and a decoder-only model to create a sophisticated encoder-decoder structure, achieving strong performance in diverse tasks [46, 110, 128], besides using a simple Transformer [29, 135] in pre-training. The variety of architectural choices underscores the potential and necessity for further exploration in this field.
The Transformer-based backbone of pre-trained ECG-FMs typically includes seven categories: ViT-series, BERT-series, CLIP-series, GPT-series, LLaMA-series, and DAll-E-series.
  • ViT-series models. ViT was originally developed to revolutionize image recognition by directly applying the Transformer architecture to sequences of image patches [20]. Unlike CNNs that rely on local feature extraction through hierarchical layers, ViT takes a unique approach by linearly embedding images into non-overlapping vectors and treating them as input tokens for the Transformer encoder. The multi-head self-attention module in ViT empowers it to capture long-range dependencies across the entire image, setting a new standard for image analysis. ViT comes in two main variants: ViT-Base and ViT-Large. ViT-Large, with its 24 layers, 1024 hidden dimensions per layer, and 16 self-attention heads, represents a leap in architectural complexity. This deeper architecture equips ViT-Large with unparalleled capacity to capture complex patterns and long-range dependencies, making it particularly effective for handling large-scale datasets. In the realm of ECG-FM, ViT has emerged as a pre-trained model, enabling the extraction of spatial and temporal cardiac features from ECG plots or time-frequency images [80, 130]. Furthermore, ViT has also been leveraged in several studies to process ECG sequences by segmenting signals into non-overlapping time series windows [41, 96, 1, 123].
@@
    BEiT leverages a discrete variational autoencoder (dVAE) to tokenize an image into patches and discrete visual tokens, subsequently replacing randomly masked image patches with a special mask embedding. These are then fed into a ViT to learn long-range dependencies. The pre-training task of BEiT, masked image modeling (MIM), closely resembles BERT’s MLM task, aiming to predict the visual tokens of the original image based on the encoding vectors of the corrupted image. Additionally, task-specific layers are directly appended to the pre-trained encoder to fine-tune the parameters on downstream tasks.
@@
    ClinicalBERT is a specialized version of the BERT model fine-tuned on clinical texts to address the unique linguistic characteristics of medical narratives such as physician notes and discharge summaries. Unlike BioBERT, it demonstrates superior performance in named entity recognition and medical natural language inference by introducing clinical notes in the pre-training stage [5]. However, it may not be as effective on de-identification tasks, where BioBERT might outperform it due to the nature of how protected health information is masked in clinical text.
    BioLinkBERT is a pre-trained LinkBERT [126] specifically tailored for the biomedical domain, leveraging the citation links within PubMed articles. In contrast to traditional clinical BERT approaches, BioLinkBERT treats the corpus as a dynamic graph of interconnected documents, utilizing citation links to enhance the learning context. Through a dual focus on MLM and document relation prediction (DRP), BioLinkBERT captures intricate dependencies and relationships across documents, leading to superior performance in tasks requiring multi-hop reasoning and comprehension of complex inter-document relations.
@@
  • GPT-series models. GPT is an autoregressive language model that generates human-like text using the Transformer architecture, initially introduced by [88] for natural language understanding tasks. The parameters learned during unsupervised pre-training are fine-tuned for specific tasks using annotated datasets. Trained on over 40 GB of text from public websites, GPT-2 [90] has demonstrated its capability to produce coherent and contextually relevant paragraphs. However, its practical application was limited due to performance issues. To address this, a more robust model, GPT-3 [13], was developed. GPT-3 employs a similar architecture to GPT-2 but implements alternating dense and locally banded sparse attention patterns in the transformer layers. With 175 billion parameters, GPT-3 can perform tasks without extensive retraining and significantly outperforms its predecessors in zero-shot and few-shot learning scenarios. The latest iteration, GPT-4 [2], expands the input prompt to include multimodal inputs such as image and audio, addressing a limitation of GPT-3, which only accepts text inputs. In the field of ECG analysis, several studies have utilized frozen GPT models to generate potential waveform details from medical knowledge bases [128, 127]. Other practitioners have attempted to create ECG-oriented GPT models, such as ECG-GPT [46] and SignalGPT [64]. However, GPT-series models still face challenges related to transparency and interpretability, hindering their practical implementation due to the issue of hallucinations [2, 42].
@@
RNNs and LSTMs are widely acknowledged for their proficiency in modeling long-range temporal data. They excel at capturing dependencies across time steps, making them particularly suitable for analyzing ECG signals where the local information contained within one heartbeat cycle is generally related to earlier parts. Compared to Transformer-based FMs, RNN-based architectures offer unique strengths in ECG pre-training due to their high efficiency and low complexity in handling long ECG sequences. A notable example is the use of LSTM in CPC [74] to predict the latent representations of ECG sequences a fixed number of steps into the future. This emerging trend of integrating RNN-based models into the pre-training stage presents an invaluable opportunity for advancing ECG analysis.

### v2

@@ 1 Introduction
In contrast, ECG foundation models (ECG-FMs) present a more versatile solution, offering several advantages [104]. On the one hand, they can perform multiple tasks, such as arrhythmia detection, HRV analysis, and even more complex tasks like predicting heart failure or stratifying patients based on risk profiles [126]. For example, FMs trained on large ECG datasets, like the PTB-XL [111] or PhysioNet’s MIMIC-III [47] waveform database, have been shown to generalize across multiple tasks with high accuracy. These models leverage the diversity of the data, learning features that are transferable across tasks and patient populations, making them better suited for real-world clinical applications where data variability is common [94]. On the other hand, ECG foundation models can incorporate multi-modal data, integrating ECG signals with other physiological or clinical data such as photoplethysmography (PPG) [58] or electronic health records (EHRs) [16], to provide a more comprehensive picture of patient health [3].
@@
In contrast to previous surveys which concentrated on general-purpose FMs or FMs covering a wide scope of medical fields, this article focuses on the specialized domain of ECG analysis. In this article, we adopt the settings in  [65] to categorize the ECG-FMs into Transformer-based models and non-Transformer-based models. Notably, while more than half of the existing studies rely on simpler deep learning models (e.g., CNN and ResNet) rather than large-scale models (e.g., BERT and LLaMA) that have millions of parameters, these simpler models hold significant potential for further development. By increasing the scale of data and enhancing model depth, they can evolve into large frameworks. Therefore, we regard all of these models as foundation models in this article, underscoring their relevance and versatility within the field. We summarize the developmental roadmap of current ECG-FMs in Figure 1, to inspire advancements and deepen comprehension in the continuously evolving landscape of FMs for ECG analysis. In brief, our major contributions are summarized as follows.
  • Comprehensive and up-to-date review. We thoroughly review the existing literature on ECG-FMs, and provide a comprehensive and up-to-date review on FMs for ECG analysis, encompassing ECG-related datasets and benchmarks, along with a wide spectrum of tasks for ECG analysis.
@@
The rest of the article is organized as follows. Section 2 illustrates the search strategy. Section 3 describes the brief background of FMs. Section 4 presents the methodology of ECG-FMs. Section 5 discusses the existing challenges and promising directions of ECG-FMs, and Section 6 concludes the article.
@@
2 Search Strategy
All the candidate articles were collected by a comprehensive search of three databases, including PubMed, arXiv, and Google Scholar using the query conditions: ‘ECG foundation model’ $||$ ‘foundation model in ECG analysis’$||$ ‘ECG self-supervised learning’. After removing duplicated papers, we included all research papers in English. Since this review mainly focuses on methodology, we included 46 papers describing methods for foundation models in ECG analysis using deep learning algorithms.
@@ 3 Background of Foundation Models
FMs are large-scale deep-learning models trained on diverse and extensive datasets in a self-supervised or semi-supervised manner [12]. FMs generally follow the pre-training and fine-tuning paradigm, which enables the model to learn a broad range of knowledge and patterns and to leverage such prior information in various downstream tasks and applications. FMs only require considerably small-scale custom data in the fine-tuning phase, showcasing remarkable adaptability and efficiency [125]. FMs have been adopted successfully across domains such as CV, NLP, and healthcare. As the data or model size massively scales up and the available data modality dramatically increases, FMs have evolved from the initial pre-trained models like BERT to contemporary large language models (LLMs), vision foundation models (VFMs), vision-language models (VLMs), and time-series foundation models (TSFMs).
@@
Vision Foundation Models.
Inspired by the success of FMs in NLP, practitioners have delved into the realm of vision FMs, encompassing VFMs and VLMs. VFMs, pre-trained on extensive image datasets, excel in comprehending image content and extracting rich semantic information. With the escalating data scale and the growing complexity of deep neural networks, VFMs have evolved from conventional CNN-based models, such as VGG [108] and EfficientNet [114], to cutting-edge Transformer-based architectures, including ViT [21], Swin Transformer [77], Linformer [122], and DeiT [116]. Moreover, many VFMs employ SSL techniques to acquire generalizable visual representations during pre-training. A prime example is MAE [41], which focuses on the correlation between visible patches and masked regions, enabling it to handle intricate visual structures and capture high-level image semantics.
@@
In addition to VFMs, considerable research efforts have also been channeled into the development of VLMs. VLMs typically incorporate a combined encoder-decoder architecture, leveraging an LLM as the core element, alongside a visual encoder used to transform images into a format compatible with the LLM. For instance, CLIP [97] utilizes contrastive learning methods to create fused representations for images and texts, demonstrating promising capabilities in open-vocabulary image classification. Similarly, SAM [52] stands out as a VLM that delivers impressive zero-shot image segmentation performance, enabling class-agnostic segmentation based on the image and a visual prompt specifying the targeted segments. These pre-trained VLMs can then be tailored to specific tasks by presenting them with a natural language description of the task and corresponding prompts. Recent studies have showcased the promising zero-shot performance of pre-trained VLMs across various downstream vision tasks, such as image classification, object detection, and semantic segmentation [78]. Additionally, research endeavors have also explored the development of models aligning multiple paired modalities, including image-text, video-text, and video-audio [38, 32].
@@
The increasing richness of time-series data and its wide range of applications have thrust TSFMs into the spotlight as a promising direction for time series analysis [46, 65]. TSFMs aim to harness the power of the FM paradigm to develop generalized models adept at understanding and forecasting time series data across diverse domains. By leveraging large-scale time series datasets, TSFMs hold the promise of attaining superior performance across a spectrum of time series tasks, offering a unified framework that can expedite research and application development in this field. Based on the nature of time series, TSFMs can be classified into general-purpose TSFMs dealing with standard time series and domain-specific TSFMs focusing on particular tasks such as traffic forecasting and clinical diagnosis. General-purpose TSFMs predominantly utilize single or multiple data modalities to craft robust models targeting specific time series tasks, such as Voice2Series [129], TS2Vec [137], and SimMTM [20] for classification; UniTime [74] for cross-domain prediction; UniTS [29] for multi-task learning; and TimesFM [18], TimeGPT-1 [30], TTMs [24], and TSMixer [23] for zero-shot forecasting; Domain-specific TSFMs have distinct focuses and architectures tailored to the requirements of the downstream tasks. In the transportation sector, TSFMs typically employ graph structures, such as GNNs, to capture the spatio-temporal interactions within dynamic traffic systems. For instance, STGCL [75] and SPGCL [64] explore the integration of contrastive learning into spatio-temporal graph forecasting, while ST-LLM [68] combines spatio-temporal information with a partially frozen LLM to enhance traffic predictions. Additionally, TSFMs in the medical domain usually place greater emphasis on capturing the hierarchical characteristics of patient data. For example, FORMED [45] can easily capture domain knowledge and be adapted to unseen cohorts with minimum training by repurposing a pre-trained TimesFM. COMET [123] utilizes a multi-granularity contrastive learning framework to exploit data consistencies across different levels (e.g., observation-level and patient-level) in various medical time series, showcasing its potential benefits in complex disease diagnosis.
@@ 4.1 Architecture Design
As shown in Figure 2, we first investigate the architecture of ECG-FMs, including backbone models and Transformer structure, concentrating on the underlying mechanisms that determine their capabilities and the applications across different ECG-based scenarios.
@@
The original Transformer model utilizes an encoder-decoder architecture with a stack of multi-head self-attention blocks. However, the most suitable transformer mode for ECG analysis is still a subject of debate, with options including encoder-only, decoder-only, and encoder-decoder [49]. Each mode offers distinct properties and is tailored to specific scenarios. The encoder-only structure, commonly found in ViTs, BERTs, and CLIPs, processes the entire input ECG sequence at once and predicts based on the understanding of the full context, making it ideal for tasks that require interpreting the entire input data, such as cardiac disease detection [130]. On the other hand, the decoder-only mode, typically seen in GPTs, LLaMAs, and DALL-Es, showcases outstanding generalization capabilities during large-scale ECG pre-training. Its proficiency in tasks such as zero-shot inference [67, 135], ECG synthesis [16], and ECG chatbot assistant [69, 136] highlights its versatility in application. Additionally, prominent research combines an encoder-only model and a decoder-only model to create a sophisticated encoder-decoder structure, achieving strong performance in diverse tasks [50, 118, 136], besides using a simple Transformer [31, 143] in pre-training. The variety of architectural choices underscores the potential and necessity for further exploration in this field.
The Transformer-based backbone of pre-trained ECG-FMs typically includes six categories: ViT-series, BERT-series, CLIP-series, GPT-series, LLaMA-series, and DAll-E-series.
  • ViT-series models. ViT was originally developed to revolutionize image recognition by directly applying the Transformer architecture to image patches [21]. Unlike CNNs that rely on local feature extraction through hierarchical layers, ViT takes a unique approach by linearly embedding images into non-overlapping vectors and treating them as input tokens for the Transformer encoder. The multi-head self-attention module in ViT empowers it to capture long-range dependencies across the entire image, setting a new standard for image analysis. ViT comes in two main variants: ViT-Base and ViT-Large. ViT-Large, with its 24 layers, 1024 hidden dimensions per layer, and 16 self-attention heads, represents a leap in architectural complexity. This deeper architecture equips ViT-Large with unparalleled capacity to capture complex patterns and long-range dependencies, making it particularly effective for handling large-scale datasets. In the realm of ECG-FM, ViT has emerged as a pre-trained model, enabling the extraction of spatial and temporal cardiac features from ECG plots or time-frequency images [87, 138]. Furthermore, ViT has also been leveraged in several studies to process ECG sequences by segmenting signals into non-overlapping time series windows [44, 103, 1, 131].
@@
    BEiT leverages a discrete variational autoencoder (dVAE) to tokenize an image into patches and discrete visual tokens, subsequently replacing randomly masked image patches with a special mask embedding. These are then fed into a ViT to learn long-range dependencies. The pre-training task of BEiT, masked image modeling (MIM), closely resembles BERT’s MLM task, aiming to predict the visual tokens of the original image based on the encoding vectors of the corrupted image. Additionally, task-specific layers are directly appended to the pre-trained encoder to fine-tune the parameters on downstream tasks. In the realm of ECG-FM, ECG-GPT [50] adopts BEiT pre-trained on ImageNet dataset as the encoder to capture the complex visual features in ECG images. HeartBEiT [118] leverages BEiT in the masked image modeling phase to learn contextual information from tokenized representations.
@@
    ClinicalBERT is a specialized version of the BERT model fine-tuned on clinical texts to address the unique linguistic characteristics of medical narratives such as physician notes and discharge summaries. Unlike BioBERT, it demonstrates superior performance in named entity recognition and medical natural language inference by introducing clinical notes in the pre-training stage [5]. However, it may not be as effective on de-identification tasks, where BioBERT might outperform it due to the nature of how protected health information is masked in clinical text. ClinicalBERT is also widely used in the field of ECG analysis. For example, ETP [66] and METS [63] employ ClinicalBERT, pre-trained on clinical notes and bio-clinical articles, as the text-encoding branch of the CLIP framework.
    BioLinkBERT is a pre-trained LinkBERT [134] specifically tailored for the biomedical domain, leveraging the citation links within PubMed articles. In contrast to traditional clinical BERT approaches, BioLinkBERT treats the corpus as a dynamic graph of interconnected documents, utilizing citation links to enhance the learning context. Through a dual focus on MLM and document relation prediction (DRP), BioLinkBERT captures intricate dependencies and relationships across documents, leading to superior performance in tasks requiring multi-hop reasoning and comprehension of complex inter-document relations. For example, CQA-ESI [136] utilizes BioLinkBERT as the textual encoder, which effectively embeds medical terminologies owing to its efficacy in processing biomedical text data.
@@
  • GPT-series models. GPT is an autoregressive language model that generates human-like text using the Transformer architecture, initially introduced by [96] for natural language understanding tasks. The parameters learned during unsupervised pre-training are fine-tuned for specific tasks using annotated datasets. Trained on over 40 GB of text from public websites, GPT-2 [98] has demonstrated its capability to produce coherent and contextually relevant paragraphs. However, its practical application was limited due to performance issues. To address this, a more robust model, GPT-3 [13], was developed. GPT-3 employs a similar architecture to GPT-2 but implements alternating dense and locally banded sparse attention patterns in the transformer layers. With 175 billion parameters, GPT-3 can perform tasks without extensive retraining and significantly outperforms its predecessors in zero-shot and few-shot learning scenarios. The latest iteration, GPT-4 [2], expands the input prompt to include multimodal inputs such as image and audio, addressing a limitation of GPT-3, which only accepts text inputs. In the field of ECG analysis, several studies have utilized frozen GPT models to generate potential waveform details from medical knowledge bases [136, 135]. Other practitioners have attempted to create ECG-oriented GPT models, such as ECG-GPT [50] and SignalGPT [69]. To mitigate diagnostic bias and improve the effectiveness of GPT fine-tuning, these specialized models first generate objective descriptions by analyzing the input multimodal ECG data, and then re-analyze the generated text using prompts resembling physicians’ corrections.
@@
RNNs and LSTMs are widely acknowledged for their proficiency in modeling long-range temporal data. They excel at capturing dependencies across time steps, making them particularly suitable for analyzing ECG signals where the local information contained within one heartbeat cycle is generally related to earlier parts. The inherent linear complexity of RNN-based architectures provides distinct advantages in efficiently processing long ECG sequences. A notable example is the use of LSTM in CPC [81]. This emerging trend of integrating RNN-based models into the pre-training stage presents an invaluable opportunity for advancing ECG analysis.